# Breaking the Global North Stereotype: A Global South-centric Benchmark Dataset for Auditing and Mitigating Biases in Facial Recognition Systems

### v1

@@ Abstract
Facial Recognition Systems (FRSs) are being developed and deployed all around the world at unprecedented rates. Most platforms are designed in a limited set of countries, but deployed in other regions too, without adequate checkpoints for region-specific requirements. This is especially problematic for Global South countries which lack strong legislation to safeguard persons facing disparate performance of these systems. A combination of unavailability of datasets, lack of understanding of how FRSs function and low-resource bias mitigation measures accentuate the problems at hand. In this work, we propose a self-curated face dataset composed of 6,579 unique male and female sports-persons (cricket players) from eight countries around the world. More than 50% of the dataset comprises individuals from the Global South countries and is demographically diverse. To aid adversarial audits and robust model training, we curate four adversarial variants of each image in the dataset, leading to more than 40,000 distinct images. We also use this dataset to benchmark five popular face recognition systems (FRSs), including both commercial and open-source FRSs, for the task of gender prediction (and country prediction for one of the open-source models as an example of red-teaming). Experiments on industrial FRSs reveal accuracies ranging from 98.2% (in case of Azure) to 38.1% (in case of Face++), with a large disparity between males and females in the Global South (max difference of 38.5% in case of Face++). Biases are also observed in all FRSs between females of the Global North and South (max difference of $\approx$ 50%). A Grad-CAM analysis shows that the nose, forehead and mouth are the regions of interest for one of the open-source FRSs. Based on this crucial observation, we design simple, low-resource bias mitigation solutions using few-shot and novel contrastive learning techniques that demonstrate a significant improvement in accuracy with disparity between males and females reducing from 50% to 1.5% in one of the settings. For the red-teaming experiment using the open-source Deepface model, we observe that simple fine-tuning is not very useful while contrastive learning brings steady benefits [^1] This work has been accepted for publication at AAAI/ACM AIES 2024..
@@ Introduction
Artificial Intelligence (AI) systems are being developed and deployed at an unprecedented rate around the world for various applications ranging from face recognition (Face++ 2023; Amazon 2023) to web search (Google 2023) and chatbots (OpenAI 2023). Often, the development is done in a single place but the deployment is done all over the world. For example, face recognition system AWS Rekognition (Amazon 2023) has been developed by Amazon in the USA, but is deployed in other countries for tasks like ID verification (Sur 2023) at airports, with no clear distinction between different geographical deployments. Thus, if the AI model is designed without considering the deployment context, the system may end up propagating stereotypes (Olier and Spadavecchia 2022) and biases (Buolamwini and Gebru 2018).
Face recognition systems: Face recognition is a group of classification tasks that involve detecting a face in an image followed by downstream tasks like gender/age/emotion detection and/or matching the input face against images stored in a database (face identification). Face recognition systems (FRSs) are increasingly deployed for highly sensitive applications like surveillance (Kamgar-Parsi, Lawson, and Kamgar-Parsi 2011), person re-identification (Rao, Lu, and Zhou 2019), policing (Krueckeberg and Ferris 2018), etc. This has put FRSs are under the scanner of researchers and policymakers alike (Fitzpatrick 2023; Dizikes 2023).
@@
Figure 1: Images from our FARFace dataset. The first row has images from Global North – Australia, New Zealand, England, South Africa. The second row has images from the Global South – India, Bangladesh, Pakistan and West Indies. The third row shows the average face for each region – Global North male, Global North female, Global South male and Global South female, generated by superimposing the images of individuals from each region.
@@
Benchmarking FRSs for the Global South: The first step toward evaluating the efficacy of any AI model is to benchmark it for a specific deployment scenario. In this work, we benchmark FRS models for the task of gender prediction. This is one of the most commonly deployed applications of FRSs, used for personalized recommendations (Chen et al. 2007; Dai, Ye, and Gong 2009), surveillance, assistive technologies, safety, and marketing (Ali et al. 2019). Some of these use-cases are sensitive and any discriminatory performance can have damaging impact on the targets (Buolamwini and Gebru 2018; Raji et al. 2020; Jaiswal et al. 2022). Since the FRSs are often developed in one region (geographical or socio-economic) and deployed in many, without any target-specific modifications, it is important to benchmark these models’ performance, specially in the Global South, where aggrieved individuals may have fewer options for recourse, legal or otherwise. This brings us to our first research question, (RQ1.) How do existing Facial Recognition Systems perform on the FARFace dataset which has more than half of the faces from the Global South?
@@
Explaining predictions of FRSs: As a next step, we attempted to understand how a model converges on its prediction. Note that this question is relevant only for the open-source models, where the internal operations are known. We use a state-of-the-art explainability tool for computer vision tasks – Grad-CAM (Selvaraju et al. 2017), to identify the regions of interest in the images for the Deepface FRS by analysing the activation maps. This gives us (RQ2.) Why do the misclassifications occur for different faces in the case of open source models? A suitable answer to this question could play a key role in designing solutions that can reduce the number of misclassifications and hence the disparity.
@@
Mitigating biases in FRSs for the Global South: Despite multiple reports on biases in FRSs for sensitive tasks like gender prediction, commercial vendors continue to develop and release these models for public use. As these black-box models cannot be debiased, researchers and activists have resorted to either calling for a complete ban of FRS usage (for the Future 2023) or various bias mitigation measures on open-source models. It is already known that training a deep model from scratch is extremely data and resource intensive and incurs a significant energy cost impacting the environment, hence re-training models may not be a good technique. It is beneficial to fine-tune existing pre-trained models to adapt them for a particular task and domain. As already pointed out, there is a significant lack of proper large-scale datasets from the Global South, a gap we aim to address with our work. We use low-resource, low-data and smart techniques like few-shot and contrastive learning to improve the efficiency of the Deepface FRS, with a specific focus on reducing the reported bias against individuals from the Global South. To perform this, we utilize our FARFace dataset, a more inclusive, complementary dataset to existing ones, with all its adversarial variants. Thus our final research question is (RQ3.) Can existing biases and lack of robustness in open-source FRS models be mitigated leading to an overall improved performance, using simple interventions?
@@
Our contributions: Here, we present a new large-scale, geographically diverse and adversarial face dataset FARFace and benchmark it on five face recognition systems for the gender prediction task– three commercial ones viz. Amazon AWS Rekognition (Amazon 2023), Microsoft Azure Face (Microsoft 2023) and Face++ (Face++ 2023) and, two open-source ones viz. DeepFace (Serengil and Ozpinar 2021) and Libfaceid (Umagat 2023) for both the standard condition (normal face images), as well as for various adversarial conditions – noisy filters of RGB, Greyscale, Spread, and an occlusion filter of face mask. We would like to state here that the original dataset collected by us is gender imbalanced (86% males). This is reflective of the gender imbalance in society, including sports (Singh 2023). To prevent further misuse, we will only be releasing a balanced subset (with all corresponding adversarial variants) for future research purposes, on request [^2] Dataset Request: https://forms.gle/2Nd9ntNcc71vvwEJ7. We observe that the gender prediction accuracy for females in Global South is worse consistently across the different FRSs. To further understand the reasons behind the misclassifications (for the open-source models), we use the Grad-CAM explanation tool and analyse its activation maps. Finally, we adopt two different paradigms (a) few shot learning, and (b) contrastive learning to mitigate the biases observed in the open source FRS. We show that our fine-tuned model also transfers well to other datasets from the Global North, thus being equally useful in that region. We also perform an alternative red-teaming experiment with the Deepface model to expose biases in tasks other than gender prediction– we use the Deepface model for the task of country prediction (and by extensions, ethnicity) from face images (Batsukh and Tsend 2016; AlBdairi, Xiao, and Alghaili 2020), specially for individuals from the Global South. The results from this experiment reinforces the cautionary message that such tasks, while easy to design, are unfair, biased and should not be propagated. The primary insights from our work are as follows:
@@
$\bullet$ Our extensive experiments show that the disparity in accuracy between males and females for gender prediction is lower in the Global North (maximum disparity 10% for commercial FRS– Face++ and 66.77% for open-source FRS– Libfaceid) than in the Global South (maximum disparity 38.51% for commercial FRS– Face++ and 83.35% for open-source FRS– Libfaceid).
@@
$\bullet$ There is a consistent trend of high disparity for adversarial inputs of females from Global South (similar to Jaiswal et al. (2022)), indicating a lack of model robustness. Grad-CAM analysis shows systematic regions of interest for male classification whereas they are random for females.
@@
$\bullet$ Finally, our proposed approaches to overcome the gender disparity in FRSs (for gender prediction task) by adopting few-shot and contrastive learning demonstrate improvements in female accuracy by 59% and 60% respectively.
@@
Face datasets: There exist a multitude of face datasets in literature meant for different face recognition tasks. Some are designed specifically for audit studies (Buolamwini and Gebru 2018) and could be balanced (Raji et al. 2020), be collected from volunteers (Ma, Correll, and Wittenbrink 2015; Ma, Kantner, and Wittenbrink 2020; Lakshmi et al. 2021), have multiple age (Zhang, Song, and Qi 2017; Eidinger, Enbar, and Hassner 2014; Parkhi, Vedaldi, and Zisserman 2015) and racial (Rothe, Timofte, and Van Gool 2018; Zhang, Song, and Qi 2017; Karkkainen and Joo 2021) groups. All these datasets have images mainly of individuals from the Global North, and have no adversarial images for robust model training or testing. We address both these gaps through our FARFace dataset.
Audit & bias mitigation on FRSs: Audits on commercial FRSs (Buolamwini and Gebru 2018; Raji et al. 2020; Jaiswal et al. 2022) and open-source FRSs (Dooley et al. 2022), have shown temporal persistence to biases. Adversarial audits (Jaiswal et al. 2022; Dooley et al. 2022; Majumdar et al. 2021) on existing datasets show that disparities get magnified for noisy inputs. Our audit lies at the intersection of (Jaiswal et al. 2022) and (Dooley et al. 2022; Majumdar et al. 2021) – we audit both commercial and open-source FRSs for the task of gender prediction. Similar to Majumdar et al. (2021) we also study the activation maps for the prediction outputs but for the task of gender prediction instead of landmark detection. Another prior work by Wang et al. (2022) proposes bias mitigation strategies by developing adversarial filters on input images in real time. However, our approach differs from the mentioned approach (FAAP (Wang et al. 2022)) in the following ways: (i) Their adversarial filters do not simulate any real-world situation and the authors do not disclose what type of filter is developed, hence it is difficult to compare against or audit for; (ii) Their adversarial filters conceal the gender and race information and improve performance for tasks like smile or hair color identification, whereas we explicitly audit the platforms for gender prediction, which has a large number of downstream applications ranging from marketing to surveillance. Furthermore, most existing approaches are either too complex or need significant changes to the pipeline (Gong, Liu, and Jain 2020; Wang and Deng 2020; Wang, Zhang, and Deng 2021; Conti et al. 2022; Wang et al. 2022). We, on the other hand, perform simple fine-tuning and contrastive learning to mitigate the observed biases.
@@
In this section, we describe the collection and preparation process of FARFace dataset followed by its basic demographic distribution based on gender and geographic region.
@@
To collect the dataset, we use Selenium tool to collect the images and the player metadata like country, name and gender. We perform the following steps to clean and preprocess the images:
@@
Adversarial variants: After the above ORIG set of images was prepared, we used the GIMP [^4] https://www.gimp.org/ image editor to create the following adversarial variants – RGB (with noise values 0.3 and 0.5), Spread and Greyscale. RGB and Spread have been used previously by Jaiswal et al. (2022) to create realistic adversarial inputs for auditing FRSs. RGB simulates the effect of edited images from social media and Spread simulates the effect blurry photos captured by cameras which may be exposed to natural elements like rain water. We introduce Greyscale for the first time to simulate the effect of a black and white photograph, thus removing the skin-tone of the individual in the photograph. We also use a popular tool MaskTheFace (Anwar and Raychowdhury 2020) to create masked variants of the ORIG set by applying blue N95 face masks on the images, to create an occlusion-based adversary. Example images for each of the adversarial variants are shown in Figure 2.
@@
Our downloaded set has 7,324 images, which reduces to 6,579 after cleaning. In the cleaned dataset 85.85% images are of males and 14.15% are of females. The size of the two RGB sets and the Spread set are the same as the ORIG set. The Greyscale set is slightly larger as it also includes some images which were greyscale by default. The Masked set is slightly smaller, because the masking tool did not identify faces in all images. Thus, we effectively get 40,055 images in the FARFace dataset after generating the five adversarial variants from the original set of images. The gender distribution in this final dataset is 85.98% males and 14.02% females. Some sample images are present in Figure 1.
Task description: While FRSs can be used for multiple tasks like face detection, gender/age detection and re-identification, in this paper we only perform gender detection from the input face since we have the ground truth data for only this task. Even though the player profiles are labelled with age, we do not experiment for it as there is no guarantee on the age being the same as when the photographs were clicked. Further a manual inspection reveals that a majority of the dataset is of young people (expected of sportspersons) and may cause a model to perform poorly when predicting for old people. We acknowledge that gender is a spectrum, but in this work, we consider it to be binary because all FRS models only predict either male or female as a label for the task of gender prediction (Keyes 2018). The ground truth labels for all individuals in our dataset are also only male or female. We also do not study face matching/re-identification because we only have 1 image per identity.
As an additional red-teaming task we also perform country prediction from the input face for one of the open source FRSs. The results show that such tasks are not only inaccurate, but also highly biased.
@@
$\bullet$ Bias-mitigation of Deepface: As has been observed in previous literature, and as we shall also show through our audits in the subsequent sections, all FRSs demonstrate biases against the females and more so in the Global South (faces of individuals with darker skin tones and different facial structures). We shall use two simple techniques to show that it is quite straigtforward to mitigate the observed biases in such open-source models. We choose Deepface for this as it is a deeper model and is better designed to detect more interesting facial features as opposed to Libfaceid; thus an analysis on this model is expected be more generalizable.
@@
(b) Contrastive learning: Contrastive learning has been proved to be very effective in handling adversarial inputs (Aggarwal et al. 2023). In this paper we take dual advantage of the contrastive learning setup. While on one hand, we teach the model the contrast between males and females, on the other hand we also teach it that the original and the adversarial versions of a face are the same. Based on this intuition, we propose a triplet loss (Schroff, Kalenichenko, and Philbin 2015) function to perform a novel contrastive learning based fine-tuning. The anchor points (any data point in a given class) ($x_{a}$) are chosen from each gender and country, the positive example for each anchor ($x_{a}^{+}$) is an adversarial image of the same individual, whereas the negative example ($x_{a}^{-}$) is an image of another person from the ORIG set of the opposite gender. We choose the ($x_{a}^{+}$) from RGB0.3 in this case to improve model robustness while reducing bias simultaneously and we choose the ($x_{a}^{-}$) from the opposite gender to create a stronger negative example for every anchor. We also experiment with a setup where the negative example is chosen from the same or opposite gender with a probability (Results are shared in the Supplementary section; see Table 17). The hyperparameters are – $\alpha:1e{-5}$, epochs: 40, optimizer: ADAM, $L_{triplet}:L_{BCE}$ = 0.8:0.2.
Held-out test set for all experiments: A held out test set of 480 images is created, composed of 60 images from each country, in a 2:1 ratio of males to females. Each setting is evaluated three times, and the avg. percentages are reported. We run our experiments [^7] Code will be made available upon acceptance. on an Ubuntu 18.04 LTS Intel(R) Xeon(R) Gold 6126 CPU server with NVIDIA Tesla P100 GPU (CUDA v11.4), 128 GB RAM and 48 cores.
@@
Next, looking at Fig. 3(c), we note that the trends are less systematic for female faces. Even on the ORIG set, the accuracies range from 90.12% (AWS Rekognition) to 23.83% (Libfaceid). In fact, the open-source FRSs consistently report poor accuracies (independent of the image type) with the highest being 59% by Deepface on GREY and lowest being 5.7% for Libfaceid on MASK. Interestingly the standard deviation trends reverse here, with commercial FRSs reporting a more stable performance – 4.14% (Face++) to 18.18% (Azure) as opposed to open-source FRSs which report a standard deviation of 17.52% (Deepface) and 20.72% (Libfaceid).
@@
Disparity in accuracy is higher for images from GS: Table 2 shows the average accuracy (micro avg. for all images in a given region) of gender prediction for male and female cricketers for each region in the original (ORIG) and RGB0.3 sets for all the five FRSs (results for a balanced set of males and females are in Supplementary section– Table 8) In the rest of this section, we choose the RGB filter to demonstrate our results since it has the highest adversarial effect on the model performances (see Figure 3). For the ORIG set, all five FRSs report high accuracy (min. 97.02%) of gender prediction for males in both Global North (GN) and Global South (GS) countries, with three FRSs performing marginally better for males in GN than in GS. In contrast, the accuracy of gender prediction for images of females from Global South is significantly lower than those from Global North for all FRSs consistently. This observation corroborates the observations made in prior literature (Buolamwini and Gebru 2018; Raji et al. 2020; Jaiswal et al. 2022) regarding less accurate performance of FRSs on darker female subgroups. Irrespective of the FRS, the gender prediction accuracy for females is always lower than males within the same region. The disparity observed for commercial FRSs range from nearly 3% in MSFT to nearly 10% in Face++ for the images of cricketers from Global North. However, this disparity increases to 10% in MSFT and 38.51% in Face++ respectively for images from Global South. The disparity in accuracy across gender is even more pronounced in the open-source FRSs. For images from Global North, the disparity ranges from 23.94% (for Deepface) to 66.67% (for Libfaceid). However for images from the Global South countries, this disparity goes upwards of 70% for both the open-source FRSs.
Looking at the RGB0.3 set, we see that there is a drop in accuracy for all FRSs on images from the Global North, but a slight improvement in accuracy for some FRSs on Global South images. Interestingly, for the Global North images, the minimum and maximum disparity between the males and females reduce to 2.5% and 52.41%, indicating that the accuracy for males reduces more than that for females. For the Global South images, the max disparity now becomes 87.26%. The disparity in accuracy between females of the two regions reduces for three of the FRSs.

### v2

@@ Abstract
Facial Recognition Systems (FRSs) are being developed and deployed all around the world at unprecedented rates. Most platforms are designed in a limited set of countries, but deployed in other regions too, without adequate checkpoints for region-specific requirements. This is especially problematic for Global South countries which lack strong legislation to safeguard persons facing disparate performance of these systems. A combination of unavailability of datasets, lack of understanding of how FRSs function and low-resource bias mitigation measures accentuate the problems at hand. In this work, we propose a self-curated face dataset composed of 6,579 unique male and female sports-persons (cricket players) from eight countries around the world. More than 50% of the dataset is composed of individuals from the Global South countries and is demographically diverse. To aid adversarial audits and robust model training, we curate four adversarial variants of each image in the dataset, leading to more than 40,000 distinct images. We also use this dataset to benchmark five popular facial recognition systems (FRSs), including both commercial and open-source FRSs, for the task of gender prediction (and country prediction for one of the open-source models as an example of red-teaming). Experiments on industrial FRSs reveal accuracies ranging from 98.2% (in case of Azure) to 38.1% (in case of Face++), with a large disparity between males and females in the Global South (max difference of 38.5% in case of Face++). Biases are also observed in all FRSs between females of the Global North and South (max difference of $\approx$ 50%). A Grad-CAM analysis shows that the nose, forehead and mouth are the regions of interest for one of the open-source FRSs. Based on this crucial observation, we design simple, low-resource bias mitigation solutions using few-shot and novel contrastive learning techniques that demonstrate a significant improvement in accuracy with disparity between males and females reducing from 50% to 1.5% in one of the settings. For the red-teaming experiment using the open-source Deepface model we observe that simple fine-tuning is not very useful while contrastive learning brings steady benefits. [^1] This work has been accepted for publication at AAAI/ACM AIES 2024.
@@ Introduction
Artificial Intelligence (AI) systems are being developed and deployed at an unprecedented rate around the world for various applications ranging from face recognition (Face++ 2023; Amazon 2023) to web search (Google 2023) and chatbots (OpenAI 2023). Often, the development is done in a single place but the deployment is done worldwide. For example, facial recognition system AWS Rekognition (Amazon 2023) has been developed by Amazon in the USA but is deployed in other countries for tasks like ID verification (Sur 2023) at airports, with no clear distinction between different geographical deployments. Thus, if the AI model is designed without considering the deployment context, the system may end up propagating stereotypes (Olier and Spadavecchia 2022) and biases (Buolamwini and Gebru 2018).
Facial recognition systems: Face recognition is a group of classification tasks that involve detecting a face in an image followed by downstream tasks like gender/age/emotion detection and/or matching the input face against images stored in a database (face identification). Facial recognition systems (FRSs) are increasingly deployed for highly sensitive applications like surveillance (Kamgar-Parsi, Lawson, and Kamgar-Parsi 2011), person re-identification (Rao, Lu, and Zhou 2019), policing (Krueckeberg and Ferris 2018), etc. This has put FRSs under the scanner of researchers and policymakers alike (Fitzpatrick 2023; Dizikes 2023).
@@
Figure 1: Images from our FARFace dataset. The first row has images from Global North – Australia, New Zealand, England and South Africa. The second row has images from the Global South – India, Bangladesh, Pakistan and West Indies. The third row shows the average face for each region – Global North male, Global North female, Global South male and Global South female, generated by superimposing the images of individuals from each region.
@@
Benchmarking FRSs for the Global South: The first step toward evaluating the efficacy of any AI model is to benchmark it for a specific deployment scenario. In this work, we benchmark FRS models for the task of gender prediction. This is one of the most commonly deployed applications of FRSs, used for personalized recommendations (Chen et al. 2007; Dai, Ye, and Gong 2009), surveillance, assistive technologies, safety, and marketing (Ali et al. 2019). Some of these use cases are sensitive, and any discriminatory performance can have a damaging impact on the targets (Buolamwini and Gebru 2018; Raji et al. 2020; Jaiswal et al. 2022). Since the FRSs are often developed in one region (geographical or socio-economic) and deployed in many, without any target-specific modifications, it is important to benchmark these models’ performance, especially in the Global South, where aggrieved individuals may have fewer options for recourse, legal or otherwise. This brings us to our first research question, (RQ1.) How do existing Facial Recognition Systems perform on the FARFace dataset, which has more than half of the faces from the Global South?
@@
Explaining predictions of FRSs: As a next step, we attempted to understand how a model converges on its prediction. Note that this question is relevant only for the open-source models, where the internal operations are known. We use a state-of-the-art explainability tool for computer vision tasks – Grad-CAM (Selvaraju et al. 2017), to identify the regions of interest in the images for the Deepface FRS by analysing the activation maps. This gives us (RQ2.) Why do the misclassifications occur for different faces in the case of open-source models? A suitable answer to this question could play a key role in designing solutions that can reduce the number of misclassifications and, hence, the disparity.
@@
Mitigating biases in FRSs for the Global South: Despite multiple reports on biases in FRSs for sensitive tasks like gender prediction, commercial vendors continue to develop and release these models for public use. As these black-box models cannot be debiased, researchers and activists have resorted to either calling for a complete ban of FRS usage (for the Future 2023) or various bias mitigation measures on open-source models. It is already known that training a deep model from scratch is extremely data and resource intensive and incurs a significant energy cost impacting the environment, hence re-training models may not be a good technique. It is beneficial to fine-tune existing pre-trained models to adapt them for a particular task and domain. As already pointed out, there is a significant lack of proper large-scale datasets from the Global South, a gap we aim to address with our work. We use low-resource, low-data and smart techniques like few-shot and contrastive learning to improve the efficiency of the Deepface FRS, with a specific focus on reducing the reported bias against individuals from the Global South. To perform this, we utilize our FARFace dataset, a more inclusive, complementary dataset to existing ones, with all its adversarial variants. Thus, our final research question is (RQ3.) Can existing biases and lack of robustness in open-source FRS models be mitigated, leading to an overall improved performance, using simple interventions?
@@
Our contributions: Here, we present a new large-scale, geographically diverse and adversarial face dataset FARFace and benchmark it on five facial recognition systems – three commercial ones viz. Amazon AWS Rekognition (Amazon 2023), Microsoft Azure Face (Microsoft 2023) and Face++ (Face++ 2023), and two open-source ones viz. DeepFace (Serengil and Ozpinar 2021) and Libfaceid (Umagat 2023) – for the gender prediction task, for both the standard condition (normal face images), as well as for various adversarial conditions – noisy filters of RGB, Greyscale, Spread, and an occlusion filter of face mask. We observe that the gender prediction accuracy for females in the Global South is worse consistently across the different FRSs. To further understand the reasons behind the misclassifications (for the open-source models), we use the Grad-CAM explainability tool and analyse its activation maps. Finally, we adopt two different paradigms, (a) few shot learning and (b) contrastive learning, to mitigate the biases observed in the open source FRS. We show that our fine-tuned model also transfers well to other datasets from the Global North, thus being equally useful in that region. We also perform an alternative red-teaming experiment with the Deepface model to expose biases in tasks other than gender prediction– we use the Deepface model for the task of country prediction (and by extensions, ethnicity) from face images (Batsukh and Tsend 2016; AlBdairi, Xiao, and Alghaili 2020), especially for individuals from the Global South. The results from this experiment reinforce the cautionary message that such tasks, while easy to design, are unfair and biased and should not be propagated. The primary insights from our work are as follows:
@@
$\bullet$ Our extensive experiments show that the disparity in accuracy between males and females for gender prediction is lower in the Global North (maximum disparity of 10% for commercial FRS– Face++ and 66.77% for open-source FRS– Libfaceid) than in the Global South (maximum disparity 38.51% for commercial FRS– Face++ and 83.35% for open-source FRS– Libfaceid).
@@
$\bullet$ There is a consistent trend of high disparity for adversarial inputs of females from Global South (similar to Jaiswal et al. (2022)), indicating a lack of model robustness. Grad-CAM analysis shows systematic regions of interest for male classification, whereas they are random for females.
@@
$\bullet$ Finally, our proposed approaches to overcome the gender disparity in FRSs (for gender prediction task) by adopting few-shot and contrastive learning demonstrate improvements in female accuracy by 59% and 60%, respectively.
@@
Dataset availability: The FARFace dataset is available for research purposes, on request [^2] Dataset Request: https://forms.gle/2Nd9ntNcc71vvwEJ7. We would like to state here that the original dataset collected by us is gender imbalanced (86% males). This is reflective of the gender imbalance in society, including sports (Singh 2023). To prevent further misuse, we are releasing only a balanced subset (with all corresponding adversarial variants).
@@
Face datasets: There exist a multitude of face datasets in literature meant for different face recognition tasks. Some are designed specifically for audit studies (Buolamwini and Gebru 2018) and could be balanced (Raji et al. 2020), be collected from volunteers (Ma, Correll, and Wittenbrink 2015; Ma, Kantner, and Wittenbrink 2020; Lakshmi et al. 2021), have multiple age (Zhang, Song, and Qi 2017; Eidinger, Enbar, and Hassner 2014; Parkhi, Vedaldi, and Zisserman 2015) and racial (Rothe, Timofte, and Van Gool 2018; Zhang, Song, and Qi 2017; Karkkainen and Joo 2021) groups. All these datasets have images mainly of individuals from the Global North and have no adversarial images for robust model training or testing. We address both these gaps through our FARFace dataset.
Audit & bias mitigation on FRSs: Audits on commercial FRSs (Buolamwini and Gebru 2018; Raji et al. 2020; Jaiswal et al. 2022) and open-source FRSs (Dooley et al. 2022), have shown temporal persistence to biases. Adversarial audits (Jaiswal et al. 2022; Dooley et al. 2022; Majumdar et al. 2021) on existing datasets show that disparities get magnified for noisy inputs. Our audit lies at the intersection of (Jaiswal et al. 2022) and (Dooley et al. 2022; Majumdar et al. 2021) – we audit both commercial and open-source FRSs for the task of gender prediction. Similar to Majumdar et al. (2021), we also study the activation maps for the prediction outputs but for the task of gender prediction instead of landmark detection. Another prior work by Wang et al. (2022) proposes bias mitigation strategies by developing adversarial filters on input images in real time. However, our approach differs from the mentioned approach (FAAP (Wang et al. 2022)) in the following ways: (i) Their adversarial filters do not simulate any real-world situation, and the authors do not disclose what type of filter is developed, hence it is difficult to compare against or audit for; (ii) Their adversarial filters conceal the gender and race information and improve performance for tasks like smile or hair color identification, whereas we explicitly audit the platforms for gender prediction, which has a large number of downstream applications ranging from marketing to surveillance. Furthermore, most existing approaches are either too complex or need significant changes to the pipeline (Gong, Liu, and Jain 2020; Wang and Deng 2020; Wang, Zhang, and Deng 2021; Conti et al. 2022; Wang et al. 2022). We, on the other hand, perform simple fine-tuning and contrastive learning to mitigate the observed biases.
@@
In this section, we describe the collection & preparation process of the FARFace dataset followed by its basic demographic distribution based on gender and geographic region.
@@
To collect the dataset, we use the Selenium tool to collect the images and the player metadata like country, name and gender. We perform the following steps to clean and preprocess the images:
@@
Adversarial variants: After the above ORIG set of images was prepared, we used the GIMP [^4] https://www.gimp.org/ image editor to create the following adversarial variants – RGB (with noise values 0.3 and 0.5), Spread and Greyscale. RGB and Spread have been used previously by Jaiswal et al. (2022) to create realistic adversarial inputs for auditing FRSs. RGB simulates the effect of edited images from social media, and Spread simulates the effect of blurry photos captured by cameras, which may be exposed to natural elements like rainwater. We introduce Greyscale for the first time to simulate the effect of a black and white photograph, thus removing the skin-tone of the individual in the photograph. We also use a popular tool MaskTheFace (Anwar and Raychowdhury 2020) to create masked variants of the ORIG set by applying blue N95 face masks on the images, to create an occlusion-based adversary. Example images for each of the adversarial variants are shown in Figure 2.
@@
Our downloaded set has 7,324 images, which reduces to 6,579 after cleaning. In the cleaned dataset, 85.85% images are of males, and 14.15% are of females. The size of the two RGB sets and the Spread set are the same as the ORIG set. The Greyscale set is slightly larger as it also includes some images which were greyscale by default. The Masked set is slightly smaller because the masking tool did not identify faces in all images. Thus, we effectively get 40,055 images in the FARFace dataset after generating the five adversarial variants from the original set of images. The gender distribution in this final dataset is 85.98% males and 14.02% females. Some sample images are present in Figure 1.
Task description: While FRSs can be used for multiple tasks like face detection, gender/age detection and identification, in this paper, we only perform gender detection from the input face since we have the ground truth data for only this task. Even though the player profiles are labelled with age, we do not experiment for it as there is no guarantee on the age being the same as when the photographs were clicked. Further, a manual inspection reveals that a majority of the dataset is of young people (expected of sportspersons) and may cause a model to perform poorly when predicting the age for old people. We acknowledge that gender is a spectrum, but in this work, we consider it to be binary because all FRS models only predict either male or female as a label for the task of gender prediction (Keyes 2018). The ground truth labels for all individuals in our dataset are also only male or female. We also do not study face verification/identification because we only have 1 image per identity.
As an additional red-teaming task, we also perform country prediction from the input face for one of the open-source FRSs. The results show that such tasks are not only inaccurate but also highly biased.
@@
$\bullet$ Bias-mitigation of Deepface: As has been observed in previous literature, and as we shall also show through our audits in the subsequent sections, all FRSs demonstrate biases against the females and more so in the Global South (faces of individuals with darker skin tones and different facial structures). We shall use two simple techniques to show that it is quite straightforward to mitigate the observed biases in such open-source models. We choose Deepface for this as it is a deeper model and is better designed to detect more interesting facial features as opposed to Libfaceid; thus an analysis on this model is expected be more generalizable.
@@
(b) Contrastive learning: Contrastive learning has proved to be very effective in handling adversarial inputs (Aggarwal et al. 2023). In this paper we take dual advantage of the contrastive learning setup. While on one hand, we teach the model the contrast between males and females, on the other hand we also teach it that the original and the adversarial versions of a face are the same. Based on this intuition, we propose a triplet loss (Schroff, Kalenichenko, and Philbin 2015) function to perform a novel contrastive learning based fine-tuning. The anchor points (any data point in a given class) ($x_{a}$) are chosen from each gender and country, the positive example for each anchor ($x_{a}^{+}$) is an adversarial image of the same individual, whereas the negative example ($x_{a}^{-}$) is an image of another person from the ORIG set of the opposite gender. We choose the ($x_{a}^{+}$) from RGB0.3 in this case to improve model robustness while reducing bias simultaneously and we choose the ($x_{a}^{-}$) from the opposite gender to create a stronger negative example for every anchor. We also experiment with a setup where the negative example is chosen from the same or opposite gender with a probability (more results are shared in the Supplementary section). The hyperparameters are – $\alpha:1e{-5}$, epochs: 40, optimizer: ADAM, $L_{triplet}:L_{BCE}$ = 0.8:0.2.
Held-out test set for all experiments: A held out test set of 480 images is created, composed of 60 images from each country, in a 2:1 ratio of males to females. Each setting is evaluated three times, and the avg. percentages are reported. We run our experiments on an Ubuntu 18.04 LTS Intel(R) Xeon(R) Gold 6126 CPU server with NVIDIA Tesla P100 GPU (CUDA v11.4), 128 GB RAM and 48 cores.
@@
Next, looking at Fig. 3(c), we note that the trends are less systematic for female faces. Even on the ORIG set, the accuracies range from 90.12% (AWS Rekognition) to 23.83% (Libfaceid). In fact, the open-source FRSs consistently report poor accuracies (independent of the image type) with the highest being 59% by Deepface on GREY and lowest being 5.7% for Libfaceid on MASK. The standard deviation trends reverse here, with commercial FRSs reporting a more stable performance – 4.14% (Face++) to 18.18% (Azure) as opposed to open-source FRSs which report a standard deviation of 17.52% (Deepface) and 20.72% (Libfaceid).
@@
Disparity in accuracy is higher for images from GS: Table 2 shows the average accuracy (micro avg. for all images in a given region) of gender prediction for male and female cricketers for each region in the original (ORIG) and RGB0.3 sets for all the five FRSs (more results for a balanced set of males and females are shared in the Supplementary section). In the rest of this section, we choose the RGB filter to demonstrate our results since it has the highest adversarial effect on the model performances (see Figure 3). For the ORIG set, all five FRSs report high accuracy (min. 97.02%) of gender prediction for males in both Global North (GN) and Global South (GS) countries, with three FRSs performing marginally better for males in GN than in GS. In contrast, the accuracy of gender prediction for images of females from Global South is significantly lower than those from Global North for all FRSs consistently. This observation corroborates the observations made in prior literature (Buolamwini and Gebru 2018; Raji et al. 2020; Jaiswal et al. 2022) regarding less accurate performance of FRSs on darker female subgroups. Irrespective of the FRS, the gender prediction accuracy for females is always lower than males within the same region. The disparity observed for commercial FRSs ranges from nearly 3% in MSFT to nearly 10% in Face++ for the images of cricketers from Global North. However, this disparity increases to 10% in MSFT and 38.51% in Face++, respectively, for images from Global South. The disparity in accuracy across gender is even more pronounced in the open-source FRSs. For images from Global North, the disparity ranges from 23.94% (for Deepface) to 66.67% (for Libfaceid). However, for images from the Global South countries, this disparity goes upwards of 70% for both the open-source FRSs.
Looking at the RGB0.3 set, we see that there is a drop in accuracy for all FRSs on images from the Global North, but a slight improvement in accuracy for some FRSs on Global South images. Interestingly, for the Global North images, the minimum and maximum disparity between the males and females reduce to 2.5% and 52.41%, indicating that the accuracy for males reduces more than that for females. For the Global South images, the max disparity now becomes 87.26%. The disparity in accuracy between females of the two regions reduces for two of the FRSs.